# Breaking the degeneracy among regular black holes with gravitational lensing

### 21 Mar 2026 v1

@@ 1 Introduction
Black holes (BHs) represent the most profound predictions of General Relativity (GR) Poisson:2009pwt ; Misner:1973prb . Despite the remarkable success of GR theory in describing gravitational phenomena, classical black hole solutions inevitably harbor spacetime singularities where curvature invariants diverge and the classical framework ceases to be valid steinbauer2023singularity ; Hawking:1973uf . This fundamental challenge has catalyzed the development of regular black holes (RBHs) which propose a modification of the interior spacetime geometry. In regular black holes, the central singularity is replaced by a finite, non-singular core while the existence of an event horizon is preserved Bambi:2023try ; Lan:2023cvz ; Bonanno:2022rvo ; Fan:2016hvf ; culetu2013regular ; bardeen1968non ; Hayward:2005gi . In most regular black hole constructions, the central region effectively mimics a de Sitter core, ensuring that curvature scalars remain finite throughout the entire manifold. Notable and influential examples are provided by the Fan-Wang black hole horizon mapping framework Fan:2016hvf , which encompasses several well known solutions, including the Culetu culetu2013regular , Bardeen bardeen1968non and Hayward black holes Hayward:2005gi . These metrics serve as robust toy models for investigating strong field gravity without the pathological issues associated with classical singularities. Consequently, regular black holes have been extensively scrutinized over the past decade across a diverse range of theoretical and observational frameworks Kurmanov:2024hpn ; Isomura:2023oqf ; Suliyeva:2025tsp ; Karmakar:2024hng ; Maeda:2021jdc ; Chiba:2017nml ; Ma:2025dee ; Toshmatov:2018cks ; Schee:2017hof .
@@
Among the various observational probes of strong gravity, gravitational lensing is paramount for testing the spacetime geometry surrounding compact objects. The deflection of light by gravitational fields was a foundational prediction of GR and remains one of its most potent observational consequences Schneider:1992bmb ; Weinberg1972-WEIGAC ; Ghaffarnejad:2014zva . In strong gravitational lensing, the weak deflection limit provides a fundamental analytical tool. The Einstein angle measured in this limit serves as a primary observable to constrain black hole parameters Keeton:2005jd ; Gao:2024ejs . However, when photons propagate in close proximity to a black hole, the deflection angle can diverge, which leads to the formation of relativistic images created by photons that loop around the compact object multiple times darwin1959gravity ; Virbhadra:1999nm . Bozza et al. Bozza:2001xd ; Bozza:2002zj ; Bozza:2008ev ; Bozza:2010xqn pioneered the strong deflection limit formalism, wherein they showed that the deflection angle diverges logarithmically as the impact parameter approaches the critical value at the photon sphere. This analytical framework allows for the derivation of key observables, such as the angular separation and magnification ratios of relativistic images. Since its inception, the strong deflection limit has become a standard methodology for exploring modified gravity and alternative black hole candidates Gao:2021lmo ; Fu:2021fxn ; Tsukamoto:2016jzh ; Ghosh:2022mka ; Cheong:2025lwp ; Vachher:2024ezs ; Turakhonov:2024xfg ; Jha:2023qqz ; Cunha:2018acu ; Manna:2018jxb ; Pang:2018jpm ; Zhang:2018yzr ; Wang:2019cuf ; Abbas:2019olp ; Lu:2019ush ; Islam:2020xmy ; Kumar:2020sag ; Chagoya:2020bqz ; Ghosh:2020spb ; Zhang:2021yyi ; Hsieh:2021scb ; Islam:2021dyk ; Islam:2021ful ; Boero:2021afh ; Molla:2021sgw ; Islam:2022ybr ; Ramadhan:2023ogm ; Kumar:2023jgh ; Molla:2023yxn ; Xie:2024dpi ; Filho:2024isd ; Molla:2024sig ; Dong:2024alq ; Zhao:2024hep ; Vishvakarma:2024icz ; Vachher:2024ait ; Guo:2024svn ; VP:2025nwe ; Lobos:2025yxe ; Hsieh:2021rru ; Li:2025pzy ; Narayan:2019imo ; Kuang:2022xjp ; Cardoso:2008bp . The strong deflection regime provides a direct observational window into near horizon geometry. This approach is further bolstered by recent milestones, such as the horizon scale images of M87* and Sgr A* captured by the Event Horizon Telescope (EHT), which have inaugurated a new era for testing strong field gravity EventHorizonTelescope:2019dse ; EventHorizonTelescope:2022wkp . Furthermore, statistical analyses leveraging EHT data have demonstrated that deviation parameters in various black hole models can be constrained to narrow intervals through $\chi^{2}$ tests of the shadow radius (see, e.g., Kumar:2018ple ; Pantig:2022ely ; Afrin:2021wlj ; Zhang:2023rsy ).
@@
To further explore the internal structures and external signatures of regular black holes, this work provides a unified investigation into the gravitational lensing features of several representative models, namely the Culetu, Bardeen and Hayward black holes, within the generalized framework proposed by Fan and Wang Fan:2016hvf . By employing Bozza’s strong deflection limit formalism, we derive analytical expressions for lensing observables and characterize the resulting relativistic images and shadow structures. Furthermore, we leverage strong lensing data to perform a $\chi^{2}$ analysis, which allows us to impose rigorous constraints on the regularization parameter $q$. To evaluate the observational sensitivity to various regularization schemes, we also examine the shadow intensity profiles generated by both static and infalling spherical accretion flows.
@@
Beyond static imaging, the next generation Event Horizon Telescope (ngEHT)’s enhanced resolution and multi-frequency monitoring Johnson:2023ynn ; Roelofs:2022lux ; Uniyal:2025uvc are crucial for resolving higher order photon rings and accretion variations. For regular black holes, this precision breaks the “macroscopic universality” that masks their microscopic details. Specifically, observations at $345GHz$ exploit a more transparent plasma regime, which bypasses the obscuration of outer layers to probe the near horizon geometry directly. These advancements provide a definitive test to distinguish non-singular spacetimes from classical GR solutions through the detection of sub-ring structures sensitive to metric corrections.
@@
The remainder of this paper is organized as follows: Section 2 introduces the unified framework for regular black hole geometries. Section 3 presents the parameter constraints derived from both the weak deflection limit (Einstein angle) and the strong deflection regime. Section 4 discusses the inherent degeneracies observed among the three regular black hole models. Section 5 explores potential methods to break these degeneracies through multi-messenger observables. Finally, we give a conclusion in section 6.
2 The regular black hole framework
@@
where $B(r)=A(r)^{-1}$ and $C(r)=r^{2}$. A unified framework encompassing the Culetu, Bardeen and Hayward solutions was proposed by Fan and Wang Fan:2016hvf . A generalized class of Fan-Wang solutions can be written as
@@
which is algebraically equivalent to the Culetu regular black hole culetu2013regular  [^1] The case $\nu=1$ is algebraically identical to the Culetu black hole culetu2013regular . In Refs. Lutfuoglu:2026fks ; Konoplya:2025ect , this form is referred to as “ black hole solutions free of curvature singularities, sourced by dark matter halos described by galactic density profiles.” For simplicity, we denote the $\nu=1$ case as the “Culetu black hole.”. In this scenario, the dimensionless parameter $q$ is interpreted as a length scale characterizing the modification of the gravitational field near the center and smoothing the curvature singularity that would otherwise appear in the Schwarzschild solution. Consequently, the spacetime geometry interpolates smoothly between a regular inner core and the standard vacuum exterior. When $\nu=2$, the metric corresponds to the Bardeen black hole bardeen1968non , which reduces to
@@
The Bardeen solution is originally proposed as the first example of a regular black hole which can be interpreted as a self-gravitating magnetic monopole arising from nonlinear electrodynamics. In this context, the parameter $q$ is associated with the magnetic charge of the configuration. Finally, for $\nu=3$, the metric Eq. (3) reproduces the Hayward black hole Hayward:2005gi which is
@@
Figure 1: The angular radius of Einstein ring $\theta_{E}$ as functions of parameter $q$ and the $\chi^{2}$ test results for Culetu, Bardeen and Hayward black holes. In left one, the data from galaxy ESO325-G004 are $\theta_{E}^{\rm obs}=2.85^{+0.55}_{-0.25}as$ Smith:2005pq ; Gao:2024ejs . The black dashed line denotes best value. The dotted lines indicate $1\sigma$ and $2\sigma$ ($\theta_{E}=2.35as$) confidence levels, while the gray and yellow shaded regions represent corresponding uncertainties in $\theta_{E}$. To provide a clearer visualization of constrained regions, we have restricted the display range of $\theta_{E}$ by excluding its upper values. In the right one, the constraints are derived from strong gravitational lensing data. The observed shadow angular radius for M87* and Sgr A* are $21\pm 1.5\,\mu\text{as}$ and $24.35\pm 3.5\,\mu\text{as}$, respectively. The $x$-axis utilizes a logarithmic scale to capture high resolution details of the deviation parameter $q$. Given that the Schwarzschild limit ($q=0$) is the best fit result, we employ a practical lower limit of $10^{-4}$ to circumvent the singularity of the logarithmic axis at zero. This serves as a numerical proxy to represent the regime where $q\geq 0$, allowing for a continuous visualization from near Schwarzschild conditions to more extreme deviations.
3 Gravitational lensing by regular black holes
We investigate the gravitational lensing signatures of Culetu, Bardeen and Hayward regular black holes in both the weak and strong field regimes. We consider photon motion confined to the equatorial plane ($\theta=\pi/2$) of a static, spherically symmetric spacetime. For a photon with a distance of closest approach $x_{0}$, the total deflection angle is given by Bozza:2002zj
@@ 3 Gravitational lensing by regular black holes
3.1 Constraints from Einstein ring
In the weak field limit ($x_{0}\gg 1$), for a system where the source, lens and observer are nearly perfectly aligned, the Einstein ring radius $\theta_{E}$ is defined as Bozza:2008ev
@@ 3.1 Constraints from Einstein ring
where the impact parameter is given by $b=D_{L}\theta_{E}$. We derive the analytical expressions for the Einstein ring in A and utilize the observed Einstein radius of the galaxy ESO325-G004, $\theta_{E}^{\rm obs}=2.85^{+0.55}_{-0.25}\,\text{as}$ Smith:2005pq , to constrain the regularization parameter $q$ for the Culetu, Bardeen and Hayward black hole models.
@@
3.2 Constraints from strong gravitational lensing
In the strong field regime, particularly near the photon sphere, we adopt the analytical formalism developed by Refs. Bozza:2001xd ; Bozza:2002zj ; Bozza:2008ev ; Bozza:2010xqn which is briefly presented in B. The appearance of the black hole shadow is governed by the behavior of null geodesics in this strong field region. We first examine the theoretical shadow radii ($b_{m}$ and $R_{sh}$) and their corresponding angular observables ($\theta_{\infty}$ and $\theta_{d}$) to establish a comparative baseline. This approach allows us to quantify how variations in the internal structure manifest in the observable shadow profile.
3.2.1 The photon sphere radius $x_{m}$, critical impact parameter $b_{m}$ and shadow radius $R_{sh}$
The shadow boundary of a static, spherically symmetric black hole is determined by unstable photon orbits. The photon sphere radius $x_{m}$ is the largest root of Bozza:2002zj
@@ 3.2.1 The photon sphere radius $x_{m}$, critical impact parameter $b_{m}$ and shadow radius $R_{sh}$
For a distant observer, the geometric shadow radius $R_{sh}$ in the celestial plane $(X,Y)$ is numerically identical to the critical impact parameter cunningham1972
@@
3.2.2 The asymptotic angular position $\theta_{\infty}$ and shadow angular radius $\theta_{d}$
@@
The angular radius of shadow, as typically reported by the EHT, is given by Abdujabbarov:2016hnw
@@
3.2.3 Constraining results by strong gravitational lensing
To constrain the parameter $q$, we perform a $\chi^{2}$ analysis based on observables sensitive to the near horizon geometry Zhang:2023rsy
@@ 3.2.3 Constraining results by strong gravitational lensing
3.3 Joint lensing analysis
@@ 4 The model degeneracy
4.1 Photon sphere radius $x_{m}$ and critical impact parameter $b_{m}$
@@
This pronounced degeneracy in $b_{m}$ can be attributed to several factors. Firstly, the shared Schwarzschild asymptotics (see Sec. 2) ensure that the dominant gravitational contribution at these scales remains similar. Secondly, $b_{m}$ is a composite quantity involving both $x_{m}$ and the metric function $A(x_{m})$. The model dependent variations in $x_{m}$ are effectively canceled out by corresponding changes in $A(x_{m})$ through a compensation mechanism that suppresses inter model differences. Finally, current observational uncertainties further obscure these subtle deviations, which renders the $b_{m}$ dominated predictions effectively indistinguishable at the $2\sigma$ confidence level.
4.2 Strong lensing observables related to $b_{m}$
Following Bozza’s analytical formalism Bozza:2002zj , we first evaluate the leading order strong lensing observables, which are essentially dictated by the critical impact parameter $b_{m}$. This shared dependence results in a pronounced geometric and temporal degeneracy, which renders different regular black hole candidates nearly indistinguishable under current observational precision.
@@ 4.2 Strong lensing observables related to $b_{m}$
4.3 Summary of existing constraints
@@
5 Potential avenues to break model degeneracy

### 26 Mar 2026 v2

@@ I Introduction
Black holes (BHs) represent the most profound predictions of General Relativity (GR) [1, 2, 3]. Despite the remarkable success of GR theory in describing gravitational phenomena, classical black hole solutions inevitably harbor spacetime singularities where curvature invariants diverge and the classical framework ceases to be valid [4, 5]. This fundamental challenge has catalyzed the development of regular black holes (RBHs) which propose a modification of the interior spacetime geometry. In regular black holes, the central singularity is replaced by a finite, non-singular core while the existence of an event horizon is preserved [6, 7, 8, 9, 10, 11, 12, 13]. In most regular black hole constructions, the central region effectively mimics a de Sitter core, ensuring that curvature scalars remain finite throughout the entire manifold. Notable and influential examples are provided by the Fan-Wang black hole horizon mapping framework [10], which encompasses several well known solutions, including the Culetu [11], Bardeen [12] and Hayward black holes [13]. These metrics serve as robust toy models for investigating strong field gravity without the pathological issues associated with classical singularities. Consequently, regular black holes have been extensively scrutinized over the past decade across a diverse range of theoretical and observational frameworks [14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
@@
Among the various observational probes of strong gravity, gravitational lensing is paramount for testing the spacetime geometry surrounding compact objects. The deflection of light by gravitational fields was a foundational prediction of GR and remains one of its most potent observational consequences [24, 25, 26]. In strong gravitational lensing, the weak deflection limit provides a fundamental analytical tool. The Einstein angle measured in this limit serves as a primary observable to constrain black hole parameters [27, 28]. However, when photons propagate in close proximity to a black hole, the deflection angle can diverge, which leads to the formation of relativistic images created by photons that loop around the compact object multiple times [29, 30]. Bozza et al. [31, 32, 33, 34] pioneered the strong deflection limit formalism, wherein they showed that the deflection angle diverges logarithmically as the impact parameter approaches the critical value at the photon sphere. This analytical framework allows for the derivation of key observables, such as the angular separation and magnification ratios of relativistic images. Since its inception, the strong deflection limit has become a standard methodology for exploring modified gravity and alternative black hole candidates [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78]. The strong deflection regime provides a direct observational window into near horizon geometry. This approach is further bolstered by recent milestones, such as the horizon scale images of M87* and Sgr A* captured by the Event Horizon Telescope (EHT), which have inaugurated a new era for testing strong field gravity [79, 80]. Furthermore, statistical analyses leveraging EHT data have demonstrated that deviation parameters in various black hole models can be constrained to narrow intervals through $\chi^{2}$ tests of the shadow radius (see, e.g., [81, 82, 83, 84]).
@@
To further explore the internal structures and external signatures of regular black holes, this work provides a unified investigation into the gravitational lensing features of several representative models, namely the Culetu, Bardeen and Hayward black holes, within the generalized framework proposed by Fan and Wang [10]. By employing Bozza’s strong deflection limit formalism, we derive analytical expressions for lensing observables and characterize the resulting relativistic images and shadow structures. Furthermore, we leverage strong lensing data to perform a $\chi^{2}$ analysis, which allows us to impose rigorous constraints on the regularization parameter $q$. To evaluate the observational sensitivity to various regularization schemes, we also examine the shadow intensity profiles generated by both static and infalling spherical accretion flows.
@@
Beyond static imaging, the next generation Event Horizon Telescope (ngEHT)’s enhanced resolution and multi-frequency monitoring [85, 86, 87] are crucial for resolving higher order photon rings and accretion variations. For regular black holes, this precision breaks the “macroscopic universality” that masks their microscopic details. Specifically, observations at $345GHz$ exploit a more transparent plasma regime, which bypasses the obscuration of outer layers to probe the near horizon geometry directly. These advancements provide a definitive test to distinguish non-singular spacetimes from classical GR solutions through the detection of sub-ring structures sensitive to metric corrections.
@@
The remainder of this paper is organized as follows: Section II introduces the unified framework for regular black hole geometries. Section III presents the parameter constraints derived from both the weak deflection limit (Einstein angle) and the strong deflection regime. Section IV discusses the inherent degeneracies observed among the three regular black hole models. Section V explores potential methods to break these degeneracies through multi-messenger observables. Finally, we give a conclusion in section VI.
II The regular black hole framework
@@
where $B(r)=A(r)^{-1}$ and $C(r)=r^{2}$. A unified framework encompassing the Culetu, Bardeen and Hayward solutions was proposed by Fan and Wang [10]. A generalized class of Fan-Wang solutions can be written as
@@
which is algebraically equivalent to the Culetu regular black hole [11] [^1] The case $\nu=1$ is algebraically identical to the Culetu black hole [11]. In Refs. [93, 94], this form is referred to as “ black hole solutions free of curvature singularities, sourced by dark matter halos described by galactic density profiles.” For simplicity, we denote the $\nu=1$ case as the “Culetu black hole.”. In this scenario, the dimensionless parameter $q$ is interpreted as a length scale characterizing the modification of the gravitational field near the center and smoothing the curvature singularity that would otherwise appear in the Schwarzschild solution. Consequently, the spacetime geometry interpolates smoothly between a regular inner core and the standard vacuum exterior. When $\nu=2$, the metric corresponds to the Bardeen black hole [12], which reduces to
@@
The Bardeen solution is originally proposed as the first example of a regular black hole which can be interpreted as a self-gravitating magnetic monopole arising from nonlinear electrodynamics. In this context, the parameter $q$ is associated with the magnetic charge of the configuration. Finally, for $\nu=3$, the metric Eq. (3) reproduces the Hayward black hole [13] which is
@@
Figure 1: The angular radius of Einstein ring $\theta_{E}$ as functions of parameter $q$ and the $\chi^{2}$ test results for Culetu, Bardeen and Hayward black holes. In left one, the data from galaxy ESO325-G004 are $\theta_{E}^{\rm obs}=2.85^{+0.55}_{-0.25}as$ [89, 28]. The black dashed line denotes best value. The dotted lines indicate $1\sigma$ and $2\sigma$ ($\theta_{E}=2.35as$) confidence levels, while the gray and yellow shaded regions represent corresponding uncertainties in $\theta_{E}$. To provide a clearer visualization of constrained regions, we have restricted the display range of $\theta_{E}$ by excluding its upper values. In the right one, the constraints are derived from strong gravitational lensing data. The observed shadow angular radius for M87* and Sgr A* are $21\pm 1.5\,\mu\text{as}$ and $24.35\pm 3.5\,\mu\text{as}$, respectively. The $x$-axis utilizes a logarithmic scale to capture high resolution details of the deviation parameter $q$. Given that the Schwarzschild limit ($q=0$) is the best fit result, we employ a practical lower limit of $10^{-4}$ to circumvent the singularity of the logarithmic axis at zero. This serves as a numerical proxy to represent the regime where $q\geq 0$, allowing for a continuous visualization from near Schwarzschild conditions to more extreme deviations.
III Gravitational lensing by regular black holes
We investigate the gravitational lensing signatures of Culetu, Bardeen and Hayward regular black holes in both the weak and strong field regimes. We consider photon motion confined to the equatorial plane ($\theta=\pi/2$) of a static, spherically symmetric spacetime. For a photon with a distance of closest approach $x_{0}$, the total deflection angle is given by [32]
@@ III Gravitational lensing by regular black holes
III.1 Constraints from Einstein ring
In the weak field limit ($x_{0}\gg 1$), for a system where the source, lens and observer are nearly perfectly aligned, the Einstein ring radius $\theta_{E}$ is defined as [33]
@@ III.1 Constraints from Einstein ring
where the impact parameter is given by $b=D_{L}\theta_{E}$. We derive the analytical expressions for the Einstein ring in A and utilize the observed Einstein radius of the galaxy ESO325-G004, $\theta_{E}^{\rm obs}=2.85^{+0.55}_{-0.25}\,\text{as}$ [89], to constrain the regularization parameter $q$ for the Culetu, Bardeen and Hayward black hole models.
@@
III.2 Constraints from strong gravitational lensing
In the strong field regime, particularly near the photon sphere, we adopt the analytical formalism developed by Refs. [31, 32, 33, 34] which is briefly presented in B. The appearance of the black hole shadow is governed by the behavior of null geodesics in this strong field region. We first examine the theoretical shadow radii ($b_{m}$ and $R_{sh}$) and their corresponding angular observables ($\theta_{\infty}$ and $\theta_{d}$) to establish a comparative baseline. This approach allows us to quantify how variations in the internal structure manifest in the observable shadow profile.
III.2.1 The photon sphere radius $x_{m}$, critical impact parameter $b_{m}$ and shadow radius $R_{sh}$
The shadow boundary of a static, spherically symmetric black hole is determined by unstable photon orbits. The photon sphere radius $x_{m}$ is the largest root of [32]
@@ III.2.1 The photon sphere radius $x_{m}$, critical impact parameter $b_{m}$ and shadow radius $R_{sh}$
For a distant observer, the geometric shadow radius $R_{sh}$ in the celestial plane $(X,Y)$ is numerically identical to the critical impact parameter [90]
@@
III.2.2 The asymptotic angular position $\theta_{\infty}$ and shadow angular radius $\theta_{d}$
@@
The angular radius of shadow, as typically reported by the EHT, is given by [91]
@@
III.2.3 Constraining results by strong gravitational lensing
To constrain the parameter $q$, we perform a $\chi^{2}$ analysis based on observables sensitive to the near horizon geometry [84]
@@ III.2.3 Constraining results by strong gravitational lensing
III.3 Joint lensing analysis
@@ IV The model degeneracy
IV.1 Photon sphere radius $x_{m}$ and critical impact parameter $b_{m}$
@@
This pronounced degeneracy in $b_{m}$ can be attributed to several factors. Firstly, the shared Schwarzschild asymptotics (see Sec. II) ensure that the dominant gravitational contribution at these scales remains similar. Secondly, $b_{m}$ is a composite quantity involving both $x_{m}$ and the metric function $A(x_{m})$. The model dependent variations in $x_{m}$ are effectively canceled out by corresponding changes in $A(x_{m})$ through a compensation mechanism that suppresses inter model differences. Finally, current observational uncertainties further obscure these subtle deviations, which renders the $b_{m}$ dominated predictions effectively indistinguishable at the $2\sigma$ confidence level.
IV.2 Strong lensing observables related to $b_{m}$
Following Bozza’s analytical formalism [32], we first evaluate the leading order strong lensing observables, which are essentially dictated by the critical impact parameter $b_{m}$. This shared dependence results in a pronounced geometric and temporal degeneracy, which renders different regular black hole candidates nearly indistinguishable under current observational precision.
@@ IV.2 Strong lensing observables related to $b_{m}$
IV.3 Summary of existing constraints
@@
V Potential avenues to break model degeneracy